hartnn (hartnn):

Ifx+y=2e^θ cos ϕ,x-y=2ie^θ sin ϕ,& u is the function of x & y,then prove that (∂^2 u)/(∂θ^2 )+(∂^2 u)/(∂ϕ^2 ) =4xy (∂^2 u)/∂x∂y

OpenStudy (anonymous):

I can only thing of one thing

OpenStudy (anonymous):

Are we thinking of the same thing?

hartnn (hartnn):

my attempt

hartnn (hartnn):

probably not :P

hartnn (hartnn):

\(\Large \dfrac {∂^2 u}{∂θ^2 }+\dfrac{∂^2 u}{∂ϕ^2 } =4xy\dfrac{ ∂^2 u}{∂x∂y} \)

OpenStudy (anonymous):

this is what i see http://puu.sh/byiYF/f8b94386cb.png

hartnn (hartnn):

refresh the page

hartnn (hartnn):

\(\Large x= e^{\theta +i\phi }\) \(\Large y= e^{\theta -i\phi }\)

OpenStudy (anonymous):

Is the first equation supposed to be \(x+y=2e^{i\theta}\) ?

hartnn (hartnn):

no just \(x+y = 2e^\theta \)

hartnn (hartnn):

i mean \(x+y = 2e^\theta \cos \phi \)

OpenStudy (anonymous):

So given \(u=f(x,y)\) and \[\large\begin{cases} x+y=2e^{\theta}\cos\varphi\\\\ x-y=2ie^{\theta}\sin\varphi \end{cases}\] you have to establish that \[\large \dfrac {∂^2 u}{∂θ^2 }+\dfrac{∂^2 u}{∂ϕ^2 } =4xy\dfrac{ ∂^2 u}{∂x∂y}\] just so I'm understanding correctly...

hartnn (hartnn):

correct

OpenStudy (anonymous):

typo: for that last equation I meant \(\dfrac{\partial^2u}{\partial\varphi^2}\)

hartnn (hartnn):

yeah, i posted that too

hartnn (hartnn):

got first partial derivate...how to get 2nd partial derivative ?

OpenStudy (anonymous):

Are you given an expression for \(u(x,y)\)?

hartnn (hartnn):

nopes

OpenStudy (anonymous):

Second partial of \(u\) w.r.t. \(\theta\). \[\begin{align*}\frac{\partial^2u}{\partial\theta^2}&=\frac{\partial}{\partial\theta}\left[\frac{\partial u}{\partial\theta}\right]\\\\ &=\frac{\partial}{\partial\theta}\left[\frac{\partial u}{\partial x}\cdot\frac{\partial x}{\partial \theta}+\frac{\partial u}{\partial y}\cdot\frac{\partial y}{\partial \theta}\right]\\\\ &=\frac{\partial }{\partial \theta}\left[\frac{\partial u}{\partial x}\cdot\frac{\partial x}{\partial \theta}\right]+\frac{\partial }{\partial \theta}\left[\frac{\partial u}{\partial y}\cdot\frac{\partial y}{\partial \theta}\right]\\\\ &=\left(\frac{\partial ^2u}{\partial \theta~\partial x }\cdot\frac{\partial x}{\partial \theta}+\frac{\partial u}{\partial x }\cdot\frac{\partial ^2x}{\partial \theta^2}\right)+\left(\frac{\partial ^2u}{\partial \theta~\partial y }\cdot\frac{\partial y}{\partial \theta}+\frac{\partial u}{\partial y }\cdot\frac{\partial ^2y}{\partial \theta^2}\right) \end{align*}\] It should be obvious that \(\dfrac{\partial x}{\partial \theta}=\dfrac{\partial x^2}{\partial \theta^2}=x\) and \(\dfrac{\partial y}{\partial \theta}=\dfrac{\partial^2 y}{\partial \theta^2}=y\), so you can reduce this a bit: \[\begin{align*} \frac{\partial^2u}{\partial\theta^2}&=\left(\frac{\partial ^2u}{\partial \theta~\partial x }\cdot\frac{\partial x}{\partial \theta}+\frac{\partial u}{\partial x }\cdot\frac{\partial ^2x}{\partial \theta^2}\right)+\left(\frac{\partial ^2u}{\partial \theta~\partial y }\cdot\frac{\partial y}{\partial \theta}+\frac{\partial u}{\partial y }\cdot\frac{\partial ^2y}{\partial \theta^2}\right)\\\\ &=x\left(\frac{\partial ^2u}{\partial \theta~\partial x }+\frac{\partial u}{\partial x }\right)+y\left(\frac{\partial ^2u}{\partial \theta~\partial y }+\frac{\partial u}{\partial y }\right)\end{align*}\]

OpenStudy (anonymous):

Doing something similar with \(\varphi\) will yield \[\begin{align*} \frac{\partial^2u}{\partial\varphi^2}&=\left(\frac{\partial ^2u}{\partial \varphi~\partial x }\cdot\frac{\partial x}{\partial \varphi}+\frac{\partial u}{\partial x }\cdot\frac{\partial ^2x}{\partial \varphi^2}\right)+\left(\frac{\partial ^2u}{\partial \varphi~\partial y }\cdot\frac{\partial y}{\partial \varphi}+\frac{\partial u}{\partial y }\cdot\frac{\partial ^2y}{\partial \varphi^2}\right)\\\\ &=\left(\frac{\partial ^2u}{\partial \varphi~\partial x }\cdot ie^{\theta+i\varphi}-\frac{\partial u}{\partial x }\cdot e^{\theta+i\varphi}\right)+\left(\frac{\partial ^2u}{\partial \varphi~\partial y }\cdot (-ie^{\theta-i\varphi})-\frac{\partial u}{\partial y }\cdot e^{\theta-i\varphi}\right)\\\\ &=x\left(i\frac{\partial ^2u}{\partial \varphi~\partial x }-\frac{\partial u}{\partial x }\right)-y\left(i\frac{\partial ^2u}{\partial \varphi~\partial y }-\frac{\partial u}{\partial y }\right)\end{align*}\]

hartnn (hartnn):

right... how will we get \(\large \dfrac{\partial^2 u}{\partial x \partial y}\) from \(\large \dfrac{\partial^2 u}{\partial \theta \partial x}\) and other such terms...

OpenStudy (anonymous):

\[\begin{align*} \frac{\partial^2u}{\partial\varphi^2} &=x\left(i\frac{\partial ^2u}{\partial \varphi~\partial x }-\frac{\partial u}{\partial x }\right)-y\left(i\frac{\partial ^2u}{\partial \varphi~\partial y }\color{red}+\frac{\partial u}{\partial y }\right)\end{align*}\]

hartnn (hartnn):

i should add them now, right ? i see few terms getting cancelled

OpenStudy (anonymous):

Not sure about that just yet... Let's see what happens when we add the partials together: \[\begin{align*} \frac{\partial^2u}{\partial\theta^2}+\frac{\partial^2u}{\partial\varphi^2}&= x\left(\frac{\partial ^2u}{\partial \theta~\partial x }+\frac{\partial u}{\partial x }\right)+y\left(\frac{\partial ^2u}{\partial \theta~\partial y }+\frac{\partial u}{\partial y }\right)\\\\ &~~~~~~~~+x\left(i\frac{\partial ^2u}{\partial \varphi~\partial x }-\frac{\partial u}{\partial x }\right)-y\left(i\frac{\partial ^2u}{\partial \varphi~\partial y }+\frac{\partial u}{\partial y }\right)\\\\ &= x\left(\frac{\partial ^2u}{\partial \theta~\partial x }+i\frac{\partial ^2u}{\partial \varphi~\partial x }\right)+y\left(\frac{\partial ^2u}{\partial \theta~\partial y }-i\frac{\partial ^2u}{\partial \varphi~\partial y }\right)\\\\ \end{align*}\]

OpenStudy (anonymous):

There's got to be a theorem relating second-order partial derivatives that lets you reduce terms like \[\frac{\partial^2u}{\partial \theta~\partial x}+\frac{\partial^2u}{\partial \theta~\partial y}~...\]

hartnn (hartnn):

\(\Large (\frac{\partial ^2u}{\partial \theta~\partial x }+i\frac{\partial ^2u}{\partial \phi~\partial x })\) should = 2y \(\frac{\partial ^2u}{\partial x~\partial y}\) and other = 2x \(\frac{\partial ^2u}{\partial x~\partial y}\) need to search some theorem that gives this result...

hartnn (hartnn):

**4 instead of 2

OpenStudy (sidsiddhartha):

hey i'm getting \[\frac{ \partial^2u }{ \partial \theta^2 }=x^2\frac{ \partial^2u }{ \partial x^2 }+2xy \frac{ \partial^2u }{ \partial y \partial x }+y^2\frac{ \partial^2u }{ \partial y^2}+x \frac{ \partial u }{ \partial x }+y \frac{ \partial u }{ \partial y }\] and \[\[\frac{ \partial^2u }{ \partial \phi^2 }=-x^2\frac{ \partial^2u }{ \partial x^2 }+2xy \frac{ \partial^2u }{ \partial y \partial x }-y^2\frac{ \partial^2u }{ \partial y^2}-x \frac{ \partial u }{ \partial x }-y \frac{ \partial u }{ \partial y }\\]\] adding them 4

hartnn (hartnn):

how ?

hartnn (hartnn):

used any theorem ?

OpenStudy (sidsiddhartha):

nope just a really lengthy calculation

hartnn (hartnn):

am i on right track ? or done something different from the beginning ?

OpenStudy (sidsiddhartha):

ure approach is correct \[\large x=e^{\theta+i \phi}\\y=e^{\theta-i \phi}\]

OpenStudy (sidsiddhartha):

now from here first \[\frac{ \partial u }{ \partial \theta }=\frac{ \partial u }{ \partial x }*\frac{ \partial x }{ \partial \theta }+\frac{ \partial u }{ \partial y }*\frac{ \partial y }{ \partial \theta }\\=e^{\theta+i \phi}*\frac{ \partial u }{ \partial x }+e^{\theta-i \phi}*\frac{ \partial u }{ \partial y }\\=x \frac{ \partial u }{ \partial x }+y \frac{ \partial u }{ \partial y }\] ok so far?

hartnn (hartnn):

yes

OpenStudy (sidsiddhartha):

now next we need to find \[\frac{ \partial^2u }{ \partial \theta^2 }\]

OpenStudy (sidsiddhartha):

\[\frac{ d^2u }{ d \theta^2 }=\frac{ d }{ d \theta }(\frac{ du }{ d \theta })\]

hartnn (hartnn):

thats what sith was doing, you're going in similar lines ?

OpenStudy (sidsiddhartha):

\[=(x \frac{ \partial }{ \partial x }+y \frac{ \partial }{ \partial y })(x \frac{ \partial u }{ \partial x }+y \frac{ \partial u }{ \partial y })\\=x^2\frac{ \partial ^2 u }{ \partial x^2 }+2xy \frac{ \partial^2u }{ \partial y \partial x }+y^2\frac{ \partial ^2u }{\partial y^2 }+x \frac{ \partial u }{ \partial x }+y \frac{ \partial u }{ \partial y }\]

OpenStudy (sidsiddhartha):

like this ok?

hartnn (hartnn):

ohh1 thats new

OpenStudy (sidsiddhartha):

and now \[\frac{ \partial u }{ \partial \phi }=\frac{ \partial u }{ \partial x }*\frac{ \partial x }{ \partial \phi }+\frac{ \partial u }{ \partial y }\frac{ \partial y }{ \partial \phi }\] same method--

OpenStudy (sidsiddhartha):

u'll get-- \[\frac{ \partial u }{ \partial \phi }=i \frac{ \partial u }{ \partial x }x-iy \frac{ \partial u }{ \partial y }\]

OpenStudy (sidsiddhartha):

again diffrentiating will give u \[\frac{ \partial^2u }{ \partial \phi^2 }\]

hartnn (hartnn):

thanks!

OpenStudy (sidsiddhartha):

no problem bohut tension mein hu campussing ke liye

hartnn (hartnn):

:O interview preparations should be started atleast 2-3 months before actual campus placements starts.... by campussing, you meant campus placements, right ?

OpenStudy (sidsiddhartha):

yeah

hartnn (hartnn):

clear concepts of C,C++ programming, data structures and algorithms, these will help you in the interview, irrespective of which branch you're in

hartnn (hartnn):

and excellent knowledge of 1 or 2 core subjects

hartnn (hartnn):

thanks to @SithsAndGiggles also, for trying :) have a look at sid's solution!

OpenStudy (anonymous):

Yup never would have guessed that... Multivariable calc...